E=180π cos108 (2πt-100y)

Find the (1) frequency in hz, (2) wave number (3) phase velocity (4) wavelength and (5) direction of propagation?


1
Expert's answer
2021-02-12T16:17:14-0500

1)

"f=\dfrac{\omega}{2\pi}=\dfrac{2\pi\ \dfrac{rad}{s}}{2\pi}=1.0\ Hz."

2)

"k=100\ m^{-1}."

3)

"v=\dfrac{\omega}{k}=\dfrac{2\pi\ \dfrac{rad}{s}}{100\ m^{-1}}=0.063\ \dfrac{m}{s}."

4)

"\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{100\ m^{-1}}=0.063\ m."

5) The wave propagates along "y"-axis.